# Overcoming Output Dimension Collapse: How Sparsity Enables Zero-shot Brain-to-Image Reconstruction at Small Data Scales

### v1

@@ 9 Acknowledgements
Appendix A Supplementary Materials
A.1 Output subspace restriction in kernel linear regression and PLS
This section shows that kernel linear regression and PLS produce predictions that are linear combinations of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$.
@@ A.1 Output subspace restriction in kernel linear regression and PLS
where $\bm{m}=(K+\lambda I)^{-1}\bm{k}_{\mathrm{te}}\in\mathbb{R}^{n}$. Thus, every prediction is a linear combination of the training outputs, independent of the particular choice of kernel.
@@
where $\bm{m}\;=\;T(T^{\top}T)^{-1}t(\bm{x}_{\mathrm{te}})\in\mathbb{R}^{n}$. Thus every prediction $\hat{\bm{y}}_{\mathrm{te}}$ is a linear combination of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$. This conclusion holds for NIPALS and SIMPLS alike, regardless of the specific construction of $T$ and $t(\cdot)$.
A.2 The best prediction
The best prediction in naive multivariate linear regression is derived analytically as follows:
@@ A.2 The best prediction
$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}},$
where $A^{+}$ is the Moore-Penrose pseudo-inverse of $A$.
@@
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$
A.3 Prediction error for sparse linear regression
@@
A.4 Selection rate-hit rate relation for correlation-based filter
@@
$\mathbf{y}_{k}$ $=X_{\mathrm{tr}}W_{:,k}+\mathbf{\epsilon}_{k}$ (31)
$=\sum_{j=1}^{d}W_{jk}\mathbf{x}_{j}+\mathbf{\epsilon}_{k},$ (32)
$=\frac{1}{\sqrt{s}}\sum_{j\in\mathcal{S}_{k}}\mathbf{x}_{j}+\mathbf{\epsilon}_{k},$ (33)
where $W_{:,k}$ is the $k$-th column of the teacher weight matrix $W$, $\mathbf{\epsilon}_{k}\in\mathbb{R}$ is the $k$-th output noise vector, and $\mathcal{S}_{k}=\left\{j\mid W_{jk}\neq 0\right\}$ is the informative index set for the $k$-th output coordinate. Within this data generation process, $\mathbf{x}_{j}\sim\mathcal{N}(\bm{0},I_{n})$ and $\mathbf{\epsilon}_{k}\sim\mathcal{N}(\bm{0},\sigma^{2}I_{n})$ hold.
@@
$=\langle\mathbf{x}_{j},\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbf{x}_{t}+\mathbf{\epsilon}_{k}\rangle$ (37)
$=\frac{1}{\sqrt{s}}\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle+\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}\setminus j}\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle+\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle.$ (38)
@@
$\mathbb{E}[z_{jk}]$ $=\frac{1}{\sqrt{s}}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle]+\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}\setminus j}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (39)
@@
$\mathrm{Var}[z_{jk}]$ $=\frac{1}{s}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle]+\frac{1}{s}\sum_{t\in\mathcal{S}_{k}\setminus j}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (41)
@@
$z_{jk}$ $=\langle\mathbf{x}_{j},\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbf{x}_{t}+\mathbf{\epsilon}_{k}\rangle$ (43)
$=\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle+\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle.$ (44)
$\mathbb{E}[z_{jk}]$ $=\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (45)
@@
$\mathrm{Var}[z_{jk}]$ $=\frac{1}{s}\sum_{t\in\mathcal{S}_{k}}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (47)
@@
(51) follows from Berry-Esseen theorem, and (53) follows from
@@
Combining (53) and (57) with (49) yields the two-equation relation
@@
A.5 Detailed Simulation Settings
We conducted simulations under the following four generative conditions:
@@ A.5 Detailed Simulation Settings
    Baseline: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{x}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$. Each non-zero entry in the teacher matrix $W$ is set to $1/\sqrt{s}$, with the remaining entries set to 0. The input vector $\bm{x}[{(i)}]$ is drawn from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
@@
    Input noise: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{z}^{(i)},\bm{x}^{(i)}=\bm{z}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$. The teacher matrix $W$ has the same binary non-zero structure as in (A). The input vector $\bm{z}^{(i)}$ is sampled from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
@@
The simulations are performed with the following parameters: the dimensionality $d=1000$ and the noise level $\sigma=0.1$.
Appendix B Supplementary Figures
Figure 1: Comparison of the best prediction and the brain prediction for subsampled training sets. Each panel shows the same analysis as in Figure 4 with subsampled training sets of 1200, 600, 300, and 150 samples. Each dot represents one image, and the dotted line indicates where the two errors are equal. Blue for natural images, orange for artificial shapes. As the data scale decreases, the output dimension collapse intensifies, and the brain’s prediction approaches the best prediction.
Figure 2: Comparison of the best prediction and the brain prediction on another reconstruction method (Ozcelik and VanRullen, 2023). (A) Left: Percentage of the best prediction error to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Blue for natural images, orange for artificial shapes. The error bars denote the standard deviation across samples. Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Each dot represents one image, and the dotted line indicates where the two errors are equal. (B) Representative reconstructions generated from the best predictions and brain predictions. From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. Similar results are observed: the brain prediction and best prediction are close to each other but far from the true feature.

### v2

@@ Acknowledgments
Appendix A Supplementary Figures
Figure A1: Comparison of the best prediction and the brain prediction for subsampled training sets. Each panel shows best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$ with subsampled training sets of 1200, 600, 300, and 150 samples. As the data scale decreases, the output dimension collapse intensifies, and the brain’s prediction approaches the best prediction.
Figure A2: Comparison of the best prediction and the brain prediction on another reconstruction method (Ozcelik and VanRullen, 2023). (A) Left: Percentage of the best prediction error to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Blue for natural images, orange for artificial shapes. The error bars denote the standard deviation across samples. Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Each dot represents one image, and the dotted line indicates where the two errors are equal. (B) Representative reconstructions generated from the best predictions and brain predictions. From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. Similar results are observed: the brain prediction and the best prediction are close to each other but far from the true feature.
Appendix B Supplementary Materials
B.1 Output subspace restriction in kernel linear regression and PLS
We show that kernel linear regression and PLS produce predictions that are linear combinations of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$. Let $(\bm{x}^{(i)},\bm{y}^{(i)})$ for $i=1,\ldots,n$ be training pairs with $\bm{x}^{(i)}\in\mathbb{R}^{d_{\mathrm{in}}}$ and $\bm{y}^{(i)}\in\mathbb{R}^{d_{\mathrm{out}}}$. Let $Y_{\mathrm{tr}}\in\mathbb{R}^{n\times d_{\mathrm{out}}}$ be the training output matrix and define $\mathcal{A}:=\{Y_{\mathrm{tr}}^{\top}\bm{m}\mid\bm{m}\in\mathbb{R}^{n}\}$.
@@ B.1 Output subspace restriction in kernel linear regression and PLS
where $\bm{m}=(K+\lambda I)^{-1}\bm{k}_{\mathrm{te}}\in\mathbb{R}^{n}$. Thus, $\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}$, i.e., every prediction is a linear combination of the training outputs, independent of the particular choice of kernel.
@@
where $\bm{m}\;=\;T(T^{\top}T)^{-1}t(\bm{x}_{\mathrm{te}})\in\mathbb{R}^{n}$. Thus, $\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}$. This conclusion holds for NIPALS and SIMPLS alike, regardless of the specific construction of $T$ and $t(\cdot)$.
B.2 Derivation of the best prediction for naive linear regression
Let $\bm{y}_{\mathrm{te}}\in\mathbb{R}^{d_{\mathrm{out}}}$ be a test output and $Y_{\mathrm{tr}}\in\mathbb{R}^{n\times d_{\mathrm{out}}}$ be the training output matrix. The best prediction in naive multivariate linear regression is derived analytically as follows:
@@ B.2 Derivation of the best prediction for naive linear regression
$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{\dagger}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}},$
where $A^{\dagger}$ is the Moore-Penrose pseudo-inverse of $A$.
@@
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{\dagger}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$
B.3 ODC analysis in the Deeprecon dataset
We analyzed output dimension collapse (ODC) on real fMRI data using the Deeprecon dataset (Shen et al., 2019). All analyses used subject sub-01 and the visual cortex (VC) ROI. Unless stated otherwise, preprocessing, latent feature extraction, and reconstruction followed Shen et al. (2019). The training set contains 1,200 ImageNet natural images (Deng et al., 2009), and the test set contains 50 natural images from held-out categories and 40 artificial shape images (Shen et al., 2019).
Our analysis mainly followed the reconstruction pipeline of Shen et al. (2019), which employs a Translator–Generator framework in which a linear translator predicts latent features from fMRI and a reconstruction step performs iterative feature inversion to produce an image that matches the target features. To analyze ODC under a naive multivariate linear translator, we replaced the sparse translator of Shen et al. (2019) with multivariate ridge regression using a fixed penalty $\lambda=1000$. Latent feature vectors were standardized using a standard scaler fit on the training outputs, and all projections and errors were computed in this standardized feature space.
We computed $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ as the orthogonal projection of the standardized true feature $\bm{y}_{\mathrm{te}}$ onto $\mathrm{span}(Y_{\mathrm{tr}})$, implemented as
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}=Y_{\mathrm{tr}}^{\top}(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top})^{\dagger}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$ (16)
To vary the data scale, we subsampled the 1,200 training images to $n\in\{600,300,150\}$ by selecting 4, 2, and 1 images per category (150 categories), respectively, and used one fixed subset for each $n$.
To visualize target features as images, we used Deep Image Prior (DIP) (Ulyanov et al., 2018) and optimized both the DIP network parameters and its input so that the features extracted from the generated image matched the target feature. We used AdamW (learning rate $10^{-3}$) for 800 iterations. Prediction error was measured as the normalized squared error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}-\bm{y}\|^{2}$. For comparing $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ and the ridge-based brain prediction $\hat{\bm{y}}_{\mathrm{te}}$, we computed the per-sample ratio $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$ and summarized it by the mean and standard deviation across samples.
In addition, we evaluated the reconstruction method of Ozcelik and VanRullen (Ozcelik and VanRullen, 2023) on the same Deeprecon data. We followed their original procedure without modification, and the corresponding results are reported in Fig. A2.
B.4 Derivation of the prediction error for sparse linear regression
@@
B.5 Correlation-based variable selection
To analyze practical sparse regression within our theoretical framework, we considered a simple correlation-based filter (akin to SIS (Fan and Lv, 2008)) as a concrete procedure (see also Fig. 7A). This choice was also motivated by practice: Shen et al. (2019) employed correlation-based voxel selection as part of their reconstruction pipeline. This method is computationally trivial compared to wrapper or embedded approaches, making it well-suited for reconstruction pipelines that must efficiently handle large-scale data. For each input vector $\mathbf{x}_{j}=\left(X_{\mathrm{tr}}\right)_{:,j}\in\mathbb{R}^{n}$ and output vector $\mathbf{y}_{k}=\left(Y_{\mathrm{tr}}\right)_{:,k}\in\mathbb{R}^{n}$ across the $n$ training samples, the model computes the sample Pearson correlation coefficient:
where $\hat{\mu}_{\mathrm{x}_{j}}=\frac{1}{n}\sum_{i=1}^{n}x_{j}^{(i)}$ and $\hat{\mu}_{\mathrm{y}_{k}}=\frac{1}{n}\sum_{i=1}^{n}y_{k}^{(i)}$ are the sample means, and $\hat{\sigma}_{\mathrm{x}_{j}}$ and $\hat{\sigma}_{\mathrm{y}_{k}}$ are the corresponding sample standard deviations. Once the model has computed the correlations $r_{jk}$ for all input-output pairs, it selects the top $d_{\mathrm{sel}}$ input variables with the highest absolute correlations $|r_{jk}|$ for each output dimension $k$. A separate ridge regression uses only the selected inputs to predict each output dimension. Applying this procedure to all $k$ yields a column-wise sparse support, which may differ across outputs.
B.6 Derivation of the selection rate and hit rate relation for correlation-based filter
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Eq. 52 follows from Berry-Esseen theorem, and Eq. 54 follows from
@@
Combining Eq. 54 and Eq. 58 with Eq. 50 yields the two-equation relation
@@
B.7 Simulation study: generative conditions and parameters
To evaluate the finite-sample accuracy of our theory and its robustness to deviations from the assumptions, we conducted simulations under the following four generative conditions:
@@ B.7 Simulation study: generative conditions and parameters
    Baseline: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{x}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$. Each non-zero entry in the teacher matrix $W$ is set to $1/\sqrt{s}$, with the remaining entries set to 0. The input vector $\bm{x}^{(i)}$ is drawn from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
@@
    Input noise: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{z}^{(i)}$ and $\bm{x}^{(i)}=\bm{z}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$. The teacher matrix $W$ has the same binary non-zero structure as in (A). The input vector $\bm{z}^{(i)}$ is sampled from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
@@
The simulations are performed with the following parameters: the dimensionality $d=1000$ and the noise level $\sigma=0.1$. For ridge regression, we used the ridge penalty that is theoretically optimal in the baseline setting, $\lambda=d/\tau^{2}$, in all simulation conditions (A–D).
B.8 Noise level estimation in the Deeprecon dataset
To set the noise level used in our theoretical comparisons with real fMRI results, we estimated the trial-to-trial noise variance from repeated presentations of identical stimuli in the Deeprecon dataset (Shen et al., 2019). We used an ANOVA-like variance decomposition to separately estimate the stimulus-driven variance and the within-stimulus trial noise.
For each voxel, let $x_{i,r}$ denote the preprocessed fMRI response to stimulus $i\in\{1,\ldots,s\}$ at repetition $r\in\{1,\ldots,t\}$. We defined the per-stimulus mean and the grand mean as
$\bar{x}_{i\cdot}=\frac{1}{t}\sum_{r=1}^{t}x_{i,r},\qquad\bar{x}_{\cdot\cdot}=\frac{1}{st}\sum_{i=1}^{s}\sum_{r=1}^{t}x_{i,r}.$ (63)
We then computed the between-stimulus and within-stimulus sums of squares,
$\mathrm{SS}_{\mathrm{between}}$ $=t\sum_{i=1}^{s}(\bar{x}_{i\cdot}-\bar{x}_{\cdot\cdot})^{2},$ (64)
$\mathrm{SS}_{\mathrm{within}}$ $=\sum_{i=1}^{s}\sum_{r=1}^{t}(x_{i,r}-\bar{x}_{i\cdot})^{2},$ (65)
and their corresponding mean squares,
$\mathrm{MS}_{\mathrm{between}}=\frac{\mathrm{SS}_{\mathrm{between}}}{s-1},\qquad\mathrm{MS}_{\mathrm{within}}=\frac{\mathrm{SS}_{\mathrm{within}}}{s(t-1)}.$ (66)
We modeled repeated responses by a random-effects decomposition
$x_{i,r}=\mu+z_{i}+\varepsilon_{i,r},$ (67)
where $z_{i}\sim\mathcal{N}(0,\sigma_{z}^{2})$ captures stimulus-driven variability across stimuli and $\varepsilon_{i,r}\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$ captures trial-to-trial noise. Under this model, the expected mean squares satisfy $\mathbb{E}[\mathrm{MS}_{\mathrm{within}}]=\sigma_{\varepsilon}^{2}$ and $\mathbb{E}[\mathrm{MS}_{\mathrm{between}}]=\sigma_{\varepsilon}^{2}+t\sigma_{z}^{2}$. We therefore estimated
$\widehat{\sigma}_{\varepsilon}^{2}=\mathrm{MS}_{\mathrm{within}},\qquad\widehat{\sigma}_{z}^{2}=\frac{\mathrm{MS}_{\mathrm{between}}-\mathrm{MS}_{\mathrm{within}}}{t}.$ (68)
We computed these estimates voxel-wise and then averaged $\widehat{\sigma}_{\varepsilon}^{2}/\widehat{\sigma}_{z}^{2}$ over voxels with $\widehat{\sigma}_{z}^{2}>0$.
Using ImageNetTraining (1200 stimuli, 5 repetitions), we obtained $\sigma_{\varepsilon}/\sigma_{z}=\sqrt{\widehat{\sigma}_{\varepsilon}^{2}/\widehat{\sigma}_{z}^{2}}\approx 5.7$. Using ImageNetTest (50 stimuli, 24 repetitions), we obtained $\sigma_{\varepsilon}/\sigma_{z}=\sqrt{\widehat{\sigma}_{\varepsilon}^{2}/\widehat{\sigma}_{z}^{2}}\approx 5.2$. Overall, these estimates suggest that the noise level $\sigma_{\varepsilon}/\sigma_{z}$ was on the order of $5$ and was relatively consistent across the two datasets.
B.9 Naive vs. sparse translator comparison in the Deeprecon dataset
We compared naive and sparse translators on the Deeprecon dataset (Shen et al., 2019) using the same subject (sub-01), VC ROI, preprocessing, and category-balanced subsampling protocol as in the ODC analysis. We evaluated prediction error on the 50 natural test images (held-out categories).
The naive translator was multivariate ridge regression with a fixed penalty $\lambda=1000$. The sparse translator was a correlation-based filter plus ridge regression: we used a fixed ridge penalty $\lambda=100$ and set the number of selected voxels per output dimension to $d_{\mathrm{sel}}\in\{63,125,250,500\}$ for $n\in\{150,300,600,1200\}$, respectively.
Prediction error was measured as the normalized squared error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}-\bm{y}\|^{2}$ and averaged across the natural test stimuli.